# Optimizing post-Newtonian parameters and fixing the BMS frame for numerical-relativity waveform hybridizations

### v1

@@ IV Hybridization procedure
In the previous section we described how we determine the PN parameters, the time offset, and the overall rotation of a PN waveform that we match to an NR waveform. However, we also need to account for differences in supertranslations and boosts between the NR and PN waveforms. While it is possible to add extra supertranslation parameters into the cost function defined in the previous section, we find it more efficient to account for all parameters (PN, time shift, boost, rotation, and supertranslations) using the following iterative procedure:
@@ V.1 Comparison between PN and NR waveforms
We present typical hybridization results for spin-aligned systems and precessing systems in Figs. 8 and 9. For the early inspiral stage, the PN and NR waveforms agree quite well, up to the systematic errors (eg., non-zero eccentricity of NR systems, absence of PN terms, etc.) that we will discuss in Sec. VI.
@@ V.2 Improvement from PN parameters and BMS frame fixing
Figure 10: Hybridization errors in the matching window, Eq. (5), for different waveforms and for different methods of choosing PN parameters and matching the BMS frame. See text for details. All the errors are measured after adjusting the 3-D rotation and time shift. The BMS frame for extrapolated waveforms cannot be modified using the method described in this paper because those waveforms do not have an accurate value of $\Psi_{2}$ to calculate the Moreschi supermomentum. We only adjust the BMS frame for CCE waveforms. Systems on the left of the vertical dotted line are spin-aligned systems; systems on the right are precessing systems, and systems are ordered left to right by $\chi_{p}$. The mass ratios for SXS:BBH:2617–2635 systems are $101<q<$ 10 and for SXS:BBH:2636–2640 systems are $q\approx 8$. The matching window is 10 orbits. For systems other than the SXS:BBH:2636–2640 systems, the matching window ends $8000\,M$ before merger. For the SXS:BBH:2636–2640 systems the matching window ends $5000\,M$ before merger because these systems are not long enough. The brown markers use the full hybridization procedure described in Section IV.
@@
In Figure 10, we present the hybridization errors of 29 systems. The hybridization error plotted is the cost function in Eq. (5) computed within a 10-orbit matching window. However, only the brown markers in the figure represent errors computed using the full hybridization procedure described in Section IV. The other markers represent results obtained using only parts of the procedure, omitting some of the waveform modes, or using NR waveforms that are not obtained using CCE; these markers are included to illustrate the need for all parts of the procedure described in Section IV. For all the markers, the 4 parameters representing the 3D rotation and time shift between the PN and NR waveforms are optimized to obtain a minimal hybridization error. The triangular markers use extrapolated NR waveforms obtained using the methods described in [42], which lack the correct memory effect. The circular markers use CCE waveforms that include memory contributions and contain enough information to allow suitable BMS transformations. The blue triangular markers denote errors obtained from the extrapolated waveforms without adding supertranslations, and with the 8 PN parameters fixed to their corresponding NR values. The extrapolated waveforms get a different CoM-correction that is described in [73]. Because the extrapolated waveforms lack the correct memory terms, we include the green triangular markers, which are obtained by setting the $m=0$ modes to zero in the NR corotating frame before hybridization, and then plotting the hybridization error of only the $m\neq 0$ modes. For the green triangular markers, the PN parameters are fixed to their corresponding NR values. The errors represented by the pink, purple, and brown circular markers include all modes including $m=0$. The pink circular markers utilize CCE waveforms, but the NR waveforms are in their inspiral superrest frames, and the PN parameters are fixed to their corresponding NR values at the reference time. The purple circular markers map the NR waveforms to the corresponding PN BMS frame, but the PN parameters used in both the PN waveform and to define the PN BMS frame are the corresponding NR values of those parameters. Comparing the pink and purple circular markers, we observe that by fixing the BMS frame we can improve the error between PN and NR waveforms by 1 to 3 orders of magnitude. This indicates that CCE NR waveforms are usually in a BMS frame that differs from the PN BMS frame, and aligning the BMS frame is crucial when comparing or hybridizing two GW waveforms.
@@
By optimizing PN parameters in addition to optimizing time shift and rotation, as shown in the red and brown, we further reduce $\mathcal{E}$ by 2 to 3 orders of magnitude for spin-aligned systems and up to 3 orders of magnitude for precessing systems. This suggests that the NR parameters at the reference time are different from the parameters that yield the best-matching PN waveform. Figure 11 shows the difference in parameters for selected systems, using the full hybridization procedure described in Section IV, which corresponds to the brown markers in Fig. 10. We see that the difference in mass ratio $q$ between PN and NR is around $10^{-2}$, the difference in total mass $M$ is around $10^{-3}\sim 10^{-2}$, and the difference in spin magnitudes is around $10^{-2}\sim 10^{-1}$.
@@
Given that the best-fit PN parameters can be significantly different from the NR parameters in some cases, it is possible that the optimization over intinsic PN parameters leads to overfitting. Missing PN terms or NR truncation errors can potentially be absorbed by this optimization, leading to an incorrect PN waveform getting attached to the NR waveform. While Figure 10 does not capture this effect because it only shows the errors in the matching window, one can check this by comparing the errors between PN and NR in an earlier test window that was not included in the optimization. This will be done in Sec. VII.
@@ VI.1 Spin-aligned systems
The blue curve in Fig. 12 shows $\mathcal{E}$ in a 15-orbit long matching window with a varying end time, and it has a local minimum when the angular velocity is around 0.012, or between 3000–8000 $M$ before merger. The increase in hybridizaton error after the minimum, where the end time of the matching window gets closer to merger, corresponds to the expected late-time PN failure that we discussed in Sec. V.1. We will show below that the increase in hybridization error before the minimum, where the matching window is moved earlier and thus farther from merger, is consistent with the effects of using an NR waveform with small but nonzero orbital eccentricity to hybridize with a PN waveform that assumes eccentricity is identically zero. The green curve in Fig. 12 will be discussed later in Sec. VI.1.
@@
Figure 13: The $L^{2}$ normalized error $\mathcal{E}$ versus eccentricity $e$ for 10 different systems: SXS:BBH:1412–1416, SXS:BBH:1153, 1164, 1165, 1168 and 3618. The end time of the matching window is $-5000\,M$, and the length of the matching window is 10 orbits. The green dotted curve is a quadratic fit to the plotted points.
@@
where $\Omega_{t}$ is the orbital frequency at time $t$, and $\Omega_{0}$ is the orbital frequency at a fiducial time when the eccentricity is $e_{0}$. According to Eqs. (11) and (12), the oscillation amplitude of the norm of the $(2,2)$ mode should be proportional to $e\propto\Omega^{-19/18}$. We check the oscillation amplitude of the norms of the $(2,2)$, $(2,-2)$, $(2,1)$ and $(2,-1)$ modes for system SXS:BBH:1412. The oscillation amplitude is obtained by subtracting the individual mode’s norm from its orbital average, which is obtained by applying a low-pass filter to the original mode. We find that the oscillation amplitudes of the norms of the above modes are proportional to $\Omega^{-19/18}$ and are thus consistent with eccentricity damping. The same argument applies to $\mathcal{E}$ within the matching window as we show in Fig. 12. Assuming that the change of $e$ and $\Omega$ in the window is small, then $\mathcal{E}\propto e^{2}\propto\Omega^{-19/9}$. The red dotted line is $e^{2}$ scaled by a coefficient found by varying the overall amplitude of the green curve until it overlays the blue curve, where the eccentricity $e$ is measured from the waveform amplitude fit of the $(2,2)$ mode using the package gw_eccentricity [74], and measurements are taken at the middle of the different matching windows. The blue curve, which is $\mathcal{E}$ between PN and NR, agrees very well with the green curve, which shows the expected effect of eccentricity damping.
@@
On the basis of these arguments, one might assume that the hybridization error would be reduced if we included eccentricity and mean anomaly in our PN parameter optimization. However, by conducting such a 14-dimensional optimization in the superrest frame of the NR system, we found that the error couldn’t be reduced further since the error generated by such a small eccentricity is overpowered by the error caused by the incorrect BMS frame. Fixing the BMS frame, however, is not a simple task when eccentricity is present, as both eccentricity and supertranslation can produce oscillations in the waveform mode amplitudes and the angular velocity. This degeneracy is depicted in Fig. 14, where we plot the eccentricity measurement in different BMS frames as a function of time. We can see by comparing the blue and orange lines that the measurement of eccentricity from the strain has strong dependence on the method of obtaining the asymptotic waveform. The measurement of eccentricity from the Bondi news—the time derivative of the strain—is less dependent on the BMS frame, but there are still large oscillations with respect to the measurement time. Because of the degeneracy between supertranslations and eccentricity, including eccentricity in hybridization is nontrivial and we leave it for future work.
@@ VI.2 Precessing systems
Figure 15: Hybridization error, Eq. (5), inside the matching window for precessing systems SXS:BBH:2617–2635. The horizontal axis is the spin-precession parameter $\chi_{p}$ for the upper panel, and $S_{\perp}/M^{2}$ for the lower panel. Here $\bm{S}_{i}=\bm{\chi}_{i}m_{i}^{2}$ is the spin angular momentum, and $S_{\perp}$ is the magnitude of the projection in the orbital plane. Results are shown for two different choices of the resolution parameter “Lev” used to compute the NR simulation. Lev4 uses a finer tolerance for the adaptive-mesh-refinement algorithm than Lev3, so Lev4 has smaller numerical errors. The matching window is 10 orbits long, and the end time is $8000\,M$ before merger.
@@
To discern whether the error is caused by the numerical error of NR spin-precession or the absence of spin-precession terms in PN, it is important to recognize that the spin terms entering the PN expansion are the spin angular momentum $\bm{S}_{i}=\bm{\chi}_{i}m_{i}^{2}$, rather than $\chi_{p}$. To investigate this, in the lower panel we plot the hybridization error for the same systems against $S_{\perp}/M^{2}=(S_{1\perp}+S_{2\perp})/M^{2}$ using two different NR simulations that are identical except for the resolution parameter “Lev” that determines the error tolerance of the simulation. Larger “Lev” means a smaller error tolerance, which translates into a more accurate (but more computationally expensive) NR simulation. By comparing the green dots and red crosses, it is evident that the NR numerical error is not the limiting error source. The slope of the green dots is $3.8$, implying that the slope of the absolute difference between PN and NR waveforms should be $1.9$. Note that the contribution of the missing PN precession-dependent spin-spin terms to the waveform is expected to be proportional to $(S_{\perp}/M^{2})^{2}$, so our result is consistent with those missing terms being the dominant source of error.
@@ VII Choice of matching window
Recall that the hybridization method involves varying the PN parameters and frame parameters to minimize the PN-NR differences in the matching window. It is not appropriate to simply compare PN to NR in the window where the difference is optimized. To better evaluate how well we match PN and NR, we also compare PN-NR differences in a new window, called a test window, that is outside the matching window so that PN-NR differences are not minimized there, and that is earlier in time than the matching window so that both PN and NR should be valid there. The idea is that for a good choice of matching window, and for sufficiently accurate NR and PN waveforms, the PN-NR differences in the test window should also be small. Unless otherwise specified, the test window is chosen to be 10 orbits long, and the center of the test window (in number of orbits) is 30 orbits earlier than the center of the matching window.
We also show the test window error for 10 long waveforms in Fig. 16. By comparing the brown and the purple columns, we can see that for most systems, the optimization over PN parameters can result in smaller error in the test window.

### v2

@@ IV Hybridization procedure
In the previous section we described how we determine the PN parameters, the time offset, and the overall rotation of a PN waveform that we match to an NR waveform. However, we also need to account for differences in supertranslations and boosts between the NR and PN waveforms. While it is possible to add extra supertranslation parameters into the cost function defined in the previous section, we find it more efficient to account for all parameters (8 PN physical parameters, time shift, boost, rotation, and supertranslations) using the following iterative procedure:
@@ V.1 Comparison between PN and NR waveforms
We present typical hybridization results for spin-aligned systems and precessing systems in Figs. 9 and 10. For the early inspiral stage, the PN and NR waveforms agree quite well, up to the systematic errors (eg., non-zero eccentricity of NR systems, absence of PN terms, etc.) that we will discuss in Sec. VI.
@@ V.2 Improvement from PN parameters and BMS frame fixing
Figure 11: Hybridization errors in the matching window, Eq. (5), for different waveforms and for different methods of choosing PN parameters and matching the BMS frame. See text for details. All the errors are measured after adjusting the 3-D rotation and time shift. The BMS frame for extrapolated waveforms cannot be modified using the method described in this paper because those waveforms do not have an accurate value of $\Psi_{2}$ to calculate the Moreschi supermomentum. We only adjust the BMS frame for CCE waveforms. Systems on the left of the vertical dotted line are spin-aligned systems; systems on the right are precessing systems, and systems are ordered left to right by $\chi_{p}$. The mass ratios for SXS:BBH:2617–2635 systems are $101<q<$ 10 and for SXS:BBH:2636–2640 systems are $q\approx 8$. The matching window is 10 orbits. For systems other than the SXS:BBH:2636–2640 systems, the matching window ends $8000\,M$ before merger. For the SXS:BBH:2636–2640 systems the matching window ends $5000\,M$ before merger because these systems are not long enough. The brown markers use the full hybridization procedure described in Section IV. $e$ is the eccentricity measured from the trajectory at the NR reference time, and $\Delta\phi$ is the dephasing between the two highest resolution NR runs of the last 20 orbits before merger. It is important to recognize that this dephasing does not directly reflect the inaccuracy of the NR waveforms, as the NR parameters within the matching window for different Levs can deviate by $10^{-2}$, given their evolution over nearly a hundred orbits prior to the matching window.
@@
Figure 13: The differences in parameters between NR and PN with respect to $\mathcal{E}$ in a 10-orbit long matching window. Each color represents a different system. The points with the same color represent the same system with different matching windows. All the systems in this figure are precessing systems.
@@
In Figure 11, we present the hybridization errors of 29 systems. The hybridization error plotted is the cost function in Eq. (5) computed within a 10-orbit matching window. However, only the brown markers in the figure represent errors computed using the full hybridization procedure described in Section IV. The other markers represent results obtained using only parts of the procedure, omitting some of the waveform modes, or using NR waveforms that are not obtained using CCE; these markers are included to illustrate the need for all parts of the procedure described in Section IV. For all the markers, the 4 parameters representing the 3D rotation and time shift between the PN and NR waveforms are optimized to obtain a minimal hybridization error. The triangular markers use extrapolated NR waveforms obtained using the methods described in [42], which lack the correct memory effect. The circular markers use CCE waveforms that include memory contributions and contain enough information to allow suitable BMS transformations. The blue triangular markers denote errors obtained from the extrapolated waveforms without adding supertranslations, and with the 8 PN parameters fixed to their corresponding NR values. The extrapolated waveforms get a different CoM-correction that is described in [76]. Because the extrapolated waveforms lack the correct memory terms, we include the green triangular markers, which are obtained by setting the $m=0$ modes to zero in the NR corotating frame before hybridization, and then plotting the hybridization error of only the $m\neq 0$ modes. For the green triangular markers, the PN parameters are fixed to their corresponding NR values. The errors represented by the pink, purple, and brown circular markers include all modes including $m=0$. The pink circular markers utilize CCE waveforms, but the NR waveforms are in their inspiral superrest frames, and the PN parameters are fixed to their corresponding NR values at the reference time. The purple circular markers map the NR waveforms to the corresponding PN BMS frame, but the PN parameters used in both the PN waveform and to define the PN BMS frame are the corresponding NR values of those parameters. Comparing the results in the superrest frame against those in the PN BMS frame (pink and purple circular markers), we observe that by fixing the BMS frame we can improve the error between PN and NR waveforms by 1 to 3 orders of magnitude. This indicates that CCE NR waveforms are usually in a BMS frame that differs from the PN BMS frame, and aligning the BMS frame is crucial when comparing or hybridizing two GW waveforms.
@@
By optimizing PN parameters in addition to optimizing time shift and rotation, as shown in the red and brown, we further reduce $\mathcal{E}$ by 2 to 3 orders of magnitude for spin-aligned systems and up to 3 orders of magnitude for precessing systems. This suggests that the NR parameters at the reference time are different from the parameters that yield the best-matching PN waveform. Figure 12 shows the difference in parameters for selected systems, using the full hybridization procedure described in Section IV, which corresponds to the brown markers in Fig. 11. We see that the fractional difference in mass ratio $q$ between PN and NR is around $10^{-2}$, the fractional difference in total mass $M$ is around $10^{-3}\sim 10^{-2}$, and the absolute difference in spin magnitudes is around $10^{-2}\sim 10^{-1}$. Fig. 13 shows the correlation between the PN/NR parameter difference and $\mathcal{E}$ in the matching window. Note that $\mathcal{E}$ is the $L^{2}$ error, so the residual in the strain is about $\mathcal{E}^{1/2}$, which is around the same order as the difference in the parameters, as expected.
@@
Given that the best-fit PN parameters can be significantly different from the NR parameters in some cases, it is possible that the optimization over intrinsic PN parameters leads to overfitting. Missing PN terms or NR truncation errors can potentially be absorbed by this optimization, leading to an incorrect PN waveform getting attached to the NR waveform. While Figure 11 does not capture this effect because it only shows the errors in the matching window, one can check this by comparing the errors between PN and NR in an earlier test window that was not included in the optimization. This will be done in Sec. VII.
@@ VI.1 Spin-aligned systems
The blue curve in Fig. 14 shows $\mathcal{E}$ in a 15-orbit long matching window with a varying end time, and it has a local minimum when the angular velocity is around 0.012, or between 3000–8000 $M$ before merger. The increase in hybridization error after the minimum, where the end time of the matching window gets closer to merger, corresponds to the expected late-time PN failure that we discussed in Sec. V.1. We will show below that the increase in hybridization error before the minimum, where the matching window is moved earlier and thus farther from merger, is consistent with the effects of using an NR waveform with small but nonzero orbital eccentricity to hybridize with a PN waveform that assumes eccentricity is identically zero. The test window error (green curve) in Fig. 14 will be discussed later in Sec. VI.1.
@@
Figure 15: The $L^{2}$ normalized error $\mathcal{E}$ versus eccentricity $e$ for 10 different systems: SXS:BBH:1412–1416, SXS:BBH:1153, 1164, 1165, 1168; and the additional system 3618. We include SXS:BBH:3618 for reference as a higher-eccentricity point ($e=0.0496$). It is spin-aligned with $\chi_{1}=\chi_{2}=0.6$ and $q=1.0$, and will be added to the public SXS catalog soon. The end time of the matching window is $-5000\,M$, and the length of the matching window is 10 orbits. The green dotted curve is a quadratic fit to the plotted points.
@@
where $\Omega_{t}$ is the orbital frequency at time $t$, and $\Omega_{0}$ is the orbital frequency at a fiducial time when the eccentricity is $e_{0}$. According to Eqs. (11) and (12), the oscillation amplitude of the norm of the $(2,2)$ mode should be proportional to $e\propto\Omega^{-19/18}$. We check the oscillation amplitude of the norms of the $(2,2)$, $(2,-2)$, $(2,1)$ and $(2,-1)$ modes for system SXS:BBH:1412. The oscillation amplitude is obtained by subtracting the individual mode’s norm from its orbital average, which is obtained by applying a low-pass filter to the original mode. We find that the oscillation amplitudes of the norms of the above modes are proportional to $\Omega^{-19/18}$ and are thus consistent with eccentricity damping. The same argument applies to $\mathcal{E}$ within the matching window as we show in Fig. 14. Assuming that the change of $e$ and $\Omega$ in the window is small, then $\mathcal{E}\propto e^{2}\propto\Omega^{-19/9}$. The red dotted line is $e^{2}$ multiplied by a coefficient chosen to approximately agree with the matching window error (blue curve), where the eccentricity $e$ is measured from the waveform amplitude fit of the $(2,2)$ mode using the package gw_eccentricity [77], and measurements are taken at the middle of the different matching windows. The value of $\mathcal{E}$ in the matching window (blue curve) agrees very well with the trend of $e^{2}$ (green curve), showing the expected effect of eccentricity damping.
@@
On the basis of these arguments, one might assume that the hybridization error would be reduced if we included eccentricity and mean anomaly in our PN parameter optimization. However, by conducting such a 14-dimensional optimization in the superrest frame of the NR system, we found that the error couldn’t be reduced further since the error generated by such a small eccentricity is overpowered by the error caused by the incorrect BMS frame. Fixing the BMS frame, however, is not a simple task when eccentricity is present, as both eccentricity and supertranslation can produce oscillations in the waveform mode amplitudes and the angular velocity. This degeneracy is depicted in Fig. 16, where we plot the eccentricity measurement in different BMS frames as a function of time. We can see by comparing the eccentricity measurement using a CCE waveform in the superrest frame (blue) and an extrapolated waveform (orange) that the measurement of eccentricity from the strain has a strong dependence on the method of obtaining the asymptotic waveform. The measurement of eccentricity from the Bondi news—the time derivative of the strain—is less dependent on the BMS frame, but there are still large oscillations with respect to the measurement time (green and red lines). Because of the degeneracy between supertranslations and eccentricity, including eccentricity in hybridization is nontrivial and we leave it for future work.
@@ VI.2 Precessing systems
Figure 17: Hybridization error, Eq. (5), inside the matching window for precessing systems SXS:BBH:2617–2635. The horizontal axis is the spin-precession parameter $\chi_{p}$ for the upper panel, and $S_{\perp}/M^{2}$ for the lower panel. Here $\bm{S}_{i}=\bm{\chi}_{i}m_{i}^{2}$ is the spin angular momentum, and $S_{\perp}$ is the magnitude of the projection in the orbital plane. Results are shown for two different choices of the resolution parameter “Lev” used to compute the NR simulation. Lev4 uses a finer tolerance for the adaptive-mesh-refinement algorithm than Lev3, so Lev4 has smaller numerical errors. The additional purple markers use only the leading order spin-spin and spin-orbit terms in both strain and $\Psi_{M}$ for Lev4, whereas the other markers use all PN terms listed in Sec. II.1, which have higher order terms in the strain, but the same terms in $\Psi_{\mathrm{M}}$. The matching window is 10 orbits long, and the end time is $8000\,M$ before merger.
@@
To discern whether the error is caused by the numerical error of NR spin-precession or the absence of spin-precession terms in PN, it is important to recognize that the spin terms entering the PN expansion are the spin angular momentum $\bm{S}_{i}=\bm{\chi}_{i}m_{i}^{2}$, rather than $\chi_{p}$. To investigate this, in the lower panel we plot the hybridization error for the same systems against $S_{\perp}/M^{2}=(S_{1\perp}+S_{2\perp})/M^{2}$ using two different NR simulations that are identical except for the resolution parameter “Lev” that determines the error tolerance of the simulation. Larger “Lev” means a smaller error tolerance, which translates into a more accurate (but more computationally expensive) NR simulation. The same Lev for all simulations used in this paper corresponds to the same tolerance for the adaptive-mesh-refinement algorithm. See the last paragraph of Sec. 2.1 in Ref. [67] for details. By comparing the results for Lev4 and Lev3 (green dots and red crosses), it is evident that the NR numerical error is not the limiting error source. The slope of the Lev4 $L^{2}$ norm error (green dots) is $3.8$, implying that the slope of the absolute difference between PN and NR waveforms should be $1.9$, since the residual is around the square root of $\mathcal{E}$. Note that the contribution of the missing PN precession-dependent spin-spin terms to the waveform is expected to be proportional to $(S_{\perp}/M^{2})^{2}$, so our result is consistent with those missing terms being the dominant source of error.
To verify consistency, we included results that only use the leading order spin-spin and spin-orbit terms in both the strain and $\Psi_{M}$ for Lev4 (purple triangular markers). It is important to recognize that this comparison is not fully consistent, as the Lev4 scenario represented by green dots also includes the spin-spin and spin-orbit PN terms at only the leading order in $\Psi_{M}$, but has up to the next-to-next leading order in the strain. Nevertheless, there are still differences when using different terms in the strain at higher values of $S_{\perp}/M^{2}$. The slope for markers including only the leading order spin terms is $3.6\pm 0.6$, suggesting that higher order PN spin terms might influence the hybridization error.
@@ VII Choice of matching window
Recall that the hybridization method involves varying the PN parameters and frame parameters to minimize the PN-NR differences in the matching window. It is not appropriate to simply compare PN to NR in the window where the difference is optimized. To better evaluate how well we match PN and NR, we also compare PN-NR differences in a new window, called a test window, that is outside the matching window so that PN-NR differences are not minimized there, and that is earlier in time than the matching window so that both PN and NR should be valid there. The idea is that for a good choice of matching window, and for sufficiently accurate NR and PN waveforms, the PN-NR differences in the test window should also be small. But it’s still expected that the differences in the test window is larger than the differences in the matching window, given that both PN and NR (waveforms and parameters) are not accurate and disagree with each other, as we explained in Sec. III and V.2. Unless otherwise specified, the test window is chosen to be 10 orbits long, and the center of the test window (in number of orbits) is 30 orbits earlier than the center of the matching window.
We also show the test window error for 10 long waveforms in Fig. 18. By comparing the results with and without optimization over PN parameters (brown and purple columns), we can see that for most systems, the optimization over PN parameters can result in smaller error in the test window.